# Towards Cross-modal Backward-compatible Representation Learning for Vision-Language Models

### v2

@@ 1 Introduction
Figure 1: A conceptual visualization of Backward-compatible Training (BT, left) and its extension, the Cross(X)-modal version (XBT, right). Circles and squares denote data samples of images and text, respectively. XBT uses Vision-Language Pretraining (VLP) models as baselines to achieve cross-modal, backward-compatible representation learning, allowing the new, improved model to be compatible with the fixed old model.
@@ 4.1 Setup
Old Model / New Model Method Text $\text{Text Query}(w,\bar{w})/{\text{Image Gallery}}(v)$ Image $\text{Image Query}(v,\bar{v})/{\text{Text Gallery}}(w)$ Case R@1 R@5 R@10 R@50 Case R@1 R@5 R@10 R@50 Original CLIP-ViT-B32 - $w/v$ 45.14 74.98 84.51 96.00 $v/w$ 71.38 92.02 96.33 99.67 CLIP-ViT-L14 - $w/v$ 47.77 76.50 85.13 95.95 $v/w$ 73.29 93.24 97.40 99.78 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Full-tune $\bar{w}_{new}/v_{old}$ 41.60 73.54 84.26 96.45 $\bar{v}_{new}/w_{old}$ 63.76 87.00 93.51 99.09 LoRA-only 43.40 74.66 84.94 96.73 68.82 90.87 96.00 99.76 Base 43.48 74.87 85.04 96.74 69.78 91.31 96.02 99.78 XBT 48.02 79.00 88.21 97.66 75.02 93.27 97.31 99.91 Full-tune $\bar{w}_{new}/\bar{v}_{new}$ 48.92 79.31 88.26 97.64 $\bar{v}_{new}/\bar{w}_{new}$ 61.22 85.89 92.62 99.04 LoRA-only 55.44 83.82 91.24 98.34 70.11 91.13 96.49 99.73 Base 56.18 84.31 91.44 98.37 71.07 91.44 96.22 99.76 XBT 63.48 88.61 93.98 98.86 77.22 94.56 97.91 99.89 CLIP-ViT-L14 / CLIP-ViT-H14 Full-tune $\bar{w}_{new}/v_{old}$ 45.18 77.82 85.34 97.65 $\bar{v}_{new}/w_{old}$ 66.38 88.55 93.59 98.90 LoRA-only 47.00 76.94 86.02 97.63 72.04 92.42 96.58 99.67 Base 47.06 77.15 86.12 97.64 73.00 92.86 96.60 99.69 XBT 51.14 80.82 88.93 98.72 76.62 94.11 97.58 99.84 Full-tune $\bar{w}_{new}/\bar{v}_{new}$ 54.39 82.26 89.06 98.95 $\bar{v}_{new}/\bar{w}_{new}$ 65.57 87.52 93.51 99.13 LoRA-only 60.91 86.77 92.04 98.65 74.46 92.73 97.38 99.82 Base 61.65 87.26 92.24 98.68 75.42 93.04 97.11 99.85 XBT 66.47 90.21 94.92 99.05 80.02 96.11 98.87 100.0
@@
[…]
@@ 4.2 Main Results
Dataset Old Model / New Model Method Text $\text{Text Query}(w,\bar{w})/{\text{Image Gallery}}(v)$ Image $\text{Image Query}(v,\bar{v})/{\text{Text Gallery}}(w)$ Case R@1 R@5 R@10 R@50 Case R@1 R@5 R@10 R@50 Large -Flickr Original CLIP-ViT-B32 - $w/v$ 21.54 41.24 50.72 72.69 $v/w$ 40.35 64.42 73.39 89.91 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Base 19.02 37.25 46.63 69.32 36.74 60.61 70.22 88.34 XBT $\bar{w}_{new}/v_{old}$ 22.38 42.71 52.41 74.80 $\bar{v}_{new}/w_{old}$ 42.47 66.04 74.87 90.79 Base 32.29 55.18 64.68 83.60 36.10 59.80 69.95 89.01 XBT $\bar{w}_{new}/\bar{v}_{new}$ 39.59 62.89 71.75 88.13 $\bar{v}_{new}/\bar{w}_{new}$ 43.50 68.50 77.99 93.21 CLIP-ViT-B32 / CLIP-ViT-H14 Base 20.27 39.17 48.59 70.67 35.53 59.44 69.06 87.67 XBT $\bar{w}_{new}/v_{old}$ 23.70 44.47 54.19 76.03 $\bar{v}_{new}/w_{old}$ 41.09 65.18 74.32 90.16 Base 32.94 55.94 65.26 83.77 34.81 59.55 69.96 88.81 XBT $\bar{w}_{new}/\bar{v}_{new}$ 40.42 63.90 72.74 88.47 $\bar{v}_{new}/\bar{w}_{new}$ 46.35 71.18 80.01 94.08 Large -COCO Original CLIP-ViT-B32 - $w/v$ 14.44 30.20 38.96 62.30 $v/w$ 28.62 50.17 59.67 80.22 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Base 12.97 28.14 36.48 59.78 26.32 46.91 56.54 77.86 XBT $\bar{w}_{new}/v_{old}$ 15.55 32.27 41.34 65.11 $\bar{v}_{new}/w_{old}$ 30.73 52.58 62.30 81.83 Base 22.16 41.65 50.98 73.06 27.26 48.16 57.80 79.57 XBT $\bar{w}_{new}/\bar{v}_{new}$ 27.61 48.69 58.08 78.83 $\bar{v}_{new}/\bar{w}_{new}$ 32.92 55.37 64.89 84.67 CLIP-ViT-B32 / CLIP-ViT-H14 Base 13.88 29.55 38.23 61.57 26.19 46.75 56.30 77.96 XBT $\bar{w}_{new}/v_{old}$ 15.82 32.79 41.84 65.45 $\bar{v}_{new}/w_{old}$ 30.33 51.28 59.97 81.66 Base 23.20 43.22 52.70 74.68 27.31 48.56 58.60 80.09 XBT $\bar{w}_{new}/\bar{v}_{new}$ 28.14 49.48 58.88 79.32 $\bar{v}_{new}/\bar{w}_{new}$ 34.54 57.29 66.79 85.92
@@
Old Model / New Model Text $\text{Text Query}(w,\bar{w})/{\text{Image Gallery}}(v)$ Image $\text{Image Query}(v,\bar{v})/{\text{Text Gallery}}(w)$ Case R@1 R@5 R@10 R@50 Case R@1 R@5 R@10 R@50 CLIP-ViT-B32 / CLIP-ViT-L14 $\bar{w}_{new}/v_{old}$ 47.68 78.00 88.00 97.50 $\bar{v}_{new}/w_{old}$ 73.89 92.00 97.22 99.86 $\bar{w}_{new}/\bar{v}_{new}$ 62.22 86.18 93.00 98.12 $\bar{v}_{new}/\bar{w}_{new}$ 76.27 92.90 96.76 99.11 CLIP-ViT-L14 / CLIP-ViT-H14 $\bar{w}_{new}/v_{old}$ 50.97 80.68 88.98 97.72 $\bar{v}_{new}/w_{old}$ 77.38 94.82 97.80 99.84 $\bar{w}_{new}/\bar{v}_{new}$ 66.09 89.78 94.77 99.01 $\bar{v}_{new}/\bar{w}_{new}$ 80.60 96.36 98.71 99.96 CLIP-ViT-B32 / CLIP-ViT-H14 $\bar{w}_{new}/v_{old}$ 49.14 79.88 88.59 97.67 $\bar{v}_{new}/w_{old}$ 74.84 93.73 97.22 99.80 $\bar{w}_{new}/\bar{v}_{new}$ 65.95 89.68 94.68 98.99 $\bar{v}_{new}/\bar{w}_{new}$ 80.96 96.00 98.71 99.96
@@
[…]
Figure 4: New query vs. Old gallery retrieval results on $nocaps$. B32 as old, and L14 as new model.
@@ 5 Conclusion
[…]

### v3

@@ 1 Introduction
Figure 1: A conceptual visualization of Backward-compatible Training (BT, above) and its extension, the Cross(X)-modal version (XBT, bottom). Circles and squares denote data samples of images and text, respectively. XBT uses Vision-Language Pretraining (VLP) models as baselines to achieve cross-modal, backward-compatible representation learning, allowing the new, improved model to be compatible with the fixed old model.
@@ 4.1 Setup
Old Model / New Model Method $\text{Text Query}(w,\bar{w})/{\text{Image Gallery}}(v)$ $\text{Image Query}(v,\bar{v})/{\text{Text Gallery}}(w)$ Case R@1 R@5 R@10 R@50 Case R@1 R@5 R@10 R@50 Original CLIP-ViT-B32 - $w/v$ 45.14 74.98 84.51 96.00 $v/w$ 71.38 92.02 96.33 99.67 CLIP-ViT-L14 - $w/v$ 47.77 76.50 85.13 95.95 $v/w$ 73.29 93.24 97.40 99.78 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Full-tune $\bar{w}_{new}/v_{old}$ 41.60 73.54 84.26 96.45 $\bar{v}_{new}/w_{old}$ 63.76 87.00 93.51 99.09 LoRA-only 43.40 74.66 84.94 96.73 68.82 90.87 96.00 99.76 Base 43.48 74.87 85.04 96.74 69.78 91.31 96.02 99.78 XBT 48.02 79.00 88.21 97.66 75.02 93.27 97.31 99.91 Full-tune $\bar{w}_{new}/\bar{v}_{new}$ 48.92 79.31 88.26 97.64 $\bar{v}_{new}/\bar{w}_{new}$ 61.22 85.89 92.62 99.04 LoRA-only 55.44 83.82 91.24 98.34 70.11 91.13 96.49 99.73 Base 56.18 84.31 91.44 98.37 71.07 91.44 96.22 99.76 XBT 63.48 88.61 93.98 98.86 77.22 94.56 97.91 99.89 CLIP-ViT-L14 / CLIP-ViT-H14 Full-tune $\bar{w}_{new}/v_{old}$ 45.18 77.82 85.34 97.65 $\bar{v}_{new}/w_{old}$ 66.38 88.55 93.59 98.90 LoRA-only 47.00 76.94 86.02 97.63 72.04 92.42 96.58 99.67 Base 47.06 77.15 86.12 97.64 73.00 92.86 96.60 99.69 XBT 51.14 80.82 88.93 98.72 76.62 94.11 97.58 99.84 Full-tune $\bar{w}_{new}/\bar{v}_{new}$ 54.39 82.26 89.06 98.95 $\bar{v}_{new}/\bar{w}_{new}$ 65.57 87.52 93.51 99.13 LoRA-only 60.91 86.77 92.04 98.65 74.46 92.73 97.38 99.82 Base 61.65 87.26 92.24 98.68 75.42 93.04 97.11 99.85 XBT 66.47 90.21 94.92 99.05 80.02 96.11 98.87 100.0
@@
In our work, we primarily use the popular CLIP models [28], based on a Transformer [36] backbone, as our baseline VLP models. The models, listed in ascending order of scale and performance, are: CLIP-ViT-B32, CLIP-ViT-L14, and CLIP-ViT-H14. For simplicity, we will refer to them as B32, L14, and H14, respectively. Notably, while B32 and L14 are trained with the same dataset at 400M scale, H14 is trained with a larger dataset at 2B scale. We apply additional LoRA [12] parameters for each new model encoder configured as follows: $\text{LoRA}_{\alpha}=16$, $\text{rank}=16$, and $\text{dropout}=0.1$. The proposed projection module $\phi$ consists of three Linear layers with layer normalization (LN) [3] and the GELU non-linearity function [9]. Detailed architecture is as follows: $\texttt{Linear}-\texttt{LN}-\texttt{GELU}-\texttt{Linear}-\texttt{LN}-\texttt{GELU}-\texttt{Linear}$. Dropout is not applied as it was found to degrade performance empirically. The intermediate hidden dimension of the Linear layer is set to be four times the dimensionality of the output embedding.
@@ 4.2 Main Results
Dataset Old Model / New Model Method $\text{Text Query}(w,\bar{w})/{\text{Image Gallery}}(v)$ $\text{Image Query}(v,\bar{v})/{\text{Text Gallery}}(w)$ Case R@1 R@5 R@10 R@50 Case R@1 R@5 R@10 R@50 Large -Flickr Original CLIP-ViT-B32 - $w/v$ 21.54 41.24 50.72 72.69 $v/w$ 40.35 64.42 73.39 89.91 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Base 19.02 37.25 46.63 69.32 36.74 60.61 70.22 88.34 XBT $\bar{w}_{new}/v_{old}$ 22.38 42.71 52.41 74.80 $\bar{v}_{new}/w_{old}$ 42.47 66.04 74.87 90.79 Base 32.29 55.18 64.68 83.60 36.10 59.80 69.95 89.01 XBT $\bar{w}_{new}/\bar{v}_{new}$ 39.59 62.89 71.75 88.13 $\bar{v}_{new}/\bar{w}_{new}$ 43.50 68.50 77.99 93.21 CLIP-ViT-B32 / CLIP-ViT-H14 Base 20.27 39.17 48.59 70.67 35.53 59.44 69.06 87.67 XBT $\bar{w}_{new}/v_{old}$ 23.70 44.47 54.19 76.03 $\bar{v}_{new}/w_{old}$ 41.09 65.18 74.32 90.16 Base 32.94 55.94 65.26 83.77 34.81 59.55 69.96 88.81 XBT $\bar{w}_{new}/\bar{v}_{new}$ 40.42 63.90 72.74 88.47 $\bar{v}_{new}/\bar{w}_{new}$ 46.35 71.18 80.01 94.08 Large -COCO Original CLIP-ViT-B32 - $w/v$ 14.44 30.20 38.96 62.30 $v/w$ 28.62 50.17 59.67 80.22 Cross-modal Backward Compatible Training CLIP-ViT-B32 / CLIP-ViT-L14 Base 12.97 28.14 36.48 59.78 26.32 46.91 56.54 77.86 XBT $\bar{w}_{new}/v_{old}$ 15.55 32.27 41.34 65.11 $\bar{v}_{new}/w_{old}$ 30.73 52.58 62.30 81.83 Base 22.16 41.65 50.98 73.06 27.26 48.16 57.80 79.57 XBT $\bar{w}_{new}/\bar{v}_{new}$ 27.61 48.69 58.08 78.83 $\bar{v}_{new}/\bar{w}_{new}$ 32.92 55.37 64.89 84.67 CLIP-ViT-B32 / CLIP-ViT-H14 Base 13.88 29.55 38.23 61.57 26.19 46.75 56.30 77.96 XBT $\bar{w}_{new}/v_{old}$ 15.82 32.79 41.84 65.45 $\bar{v}_{new}/w_{old}$ 30.33 51.28 59.97 81.66 Base 23.20 43.22 52.70 74.68 27.31 48.56 58.60 80.09 XBT $\bar{w}_{new}/\bar{v}_{new}$ 28.14 49.48 58.88 79.32 $\bar{v}_{new}/\bar{w}_{new}$ 34.54 57.29 66.79 85.92
@@
Old Model / New Model $\text{Text Query}(w,\bar{w})/{\text{Image Gallery}}(v)$ $\text{Image Query}(v,\bar{v})/{\text{Text Gallery}}(w)$ Case R@1 R@5 R@10 R@50 Case R@1 R@5 R@10 R@50 CLIP-ViT-B32 / CLIP-ViT-L14 $\bar{w}_{new}/v_{old}$ 47.68 78.00 88.00 97.50 $\bar{v}_{new}/w_{old}$ 73.89 92.00 97.22 99.86 $\bar{w}_{new}/\bar{v}_{new}$ 62.22 86.18 93.00 98.12 $\bar{v}_{new}/\bar{w}_{new}$ 76.27 92.90 96.76 99.11 CLIP-ViT-L14 / CLIP-ViT-H14 $\bar{w}_{new}/v_{old}$ 50.97 80.68 88.98 97.72 $\bar{v}_{new}/w_{old}$ 77.38 94.82 97.80 99.84 $\bar{w}_{new}/\bar{v}_{new}$ 66.09 89.78 94.77 99.01 $\bar{v}_{new}/\bar{w}_{new}$ 80.60 96.36 98.71 99.96 CLIP-ViT-B32 / CLIP-ViT-H14 $\bar{w}_{new}/v_{old}$ 49.14 79.88 88.59 97.67 $\bar{v}_{new}/w_{old}$ 74.84 93.73 97.22 99.80 $\bar{w}_{new}/\bar{v}_{new}$ 65.95 89.68 94.68 98.99 $\bar{v}_{new}/\bar{w}_{new}$ 80.96 96.00 98.71 99.96
@@
Method $\bar{w}\text{(text)}/v\text{(image)}$ $\bar{v}\text{(image)}/w\text{(text)}$ R@1 R@10 R@1 R@10 (a) Baseline 48.02 79.00 75.02 93.27 (b) w.o noise 47.24 78.05 73.31 92.49 (c) $0.25\times D_{T}$ 46.89 77.91 73.20 92.18 (d) $0.5\times D_{T}$ 47.68 78.00 73.89 92.00 (e) $2\times D$ 47.63 78.69 74.98 93.98 (f) $0.5\times D$ 46.80 77.89 71.69 92.00 (g) DD=CC3M 46.44 77.76 71.33 91.84 (h) Image-only 45.32 76.92 74.02 92.35
Figure 4: New query vs. Old gallery retrieval results on nocapsnocaps. B32 as old, and L14 as new model.
@@ 5 Conclusion
Acknowledgments
The first author performed the initial research and development of this work in collaboration with Ser-nam Lim. The finalization and refinement of this work were completed while the first author was affiliated with Google DeepMind.
@@ Acknowledgments
Method $\text{Text Query}/{\text{Image Gallery}}$ $\text{Image Query}/{\text{Text Gallery}}$ Case R@10 R@50 Case R@10 R@50 Original - $w/v$ 69.56 91.86 $v/w$ 85.44 97.89 Cross-modal Backward Compatible Training Base 67.61 91.56 81.42 97.38 XBT $\bar{w}_{new}/v_{old}$ 69.90 92.52 $\bar{v}_{new}/w_{old}$ 85.83 98.04 Base 69.02 90.57 79.29 95.98 XBT $\bar{w}_{new}/\bar{v}_{new}$ 73.17 93.27 $\bar{v}_{new}/\bar{w}_{new}$ 89.16 98.88